OpenStudy (anonymous):

A triangular prism has a volume of 64 yd3. If each dimension of the prism is reduced by half, what is the new volume of the prism?

OpenStudy (bahrom7893):

what's the formula for the volume of a triangular prism?

OpenStudy (anonymous):

V' = (1/2) V = 64/2 = 32 yd^3